There are 10 points in a plane, out of these 6 are collinear. If N is the number of triangles formed by joining these points. then

Grade:11

1 Answers

Vikas TU
3 years ago
The number of triangles formed =C(10,3)-C(6,3)
=120-20=100...................................................................................